OpenStudy (anonymous):

How do you group 2x² + 13x + 21 ?

OpenStudy (anonymous):

Given that your quadratic is in the form of ax^2 + bx +c, multiply ac to get 42. Find factors of 42 that add to 13 (the b coefficient). Substitute those coefficients for the middle term, then factor by grouping. In this case, those factors are six and seven.\[2x² + 13x + 21=2x^2+6x+7x+21\]\[=2x(x+3)+7(x+3)=(2x+7)(x+3)\]

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

Do you get it?

OpenStudy (anonymous):

Kind of. For every trinomial like this do you rewrite the middle term?

OpenStudy (anonymous):

Not every trinomial will be factorable. This method only works if the zeros of the function are rational numbers (fractions).

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

This method is usually only used for problems that have a leading coefficient (a) different than one.

OpenStudy (anonymous):

Do math every day.

OpenStudy (anonymous):

Okay thanks